# The Open Journal of Astrophysics Sulphur abundances in star-forming regions from optical emission lines: A new approach based on photoionization models consistent with the direct method

### v1

@@ Abstract
The derivation of sulphur chemical abundances in the gas-phase of star-forming galaxies is explored in this work, using the emission lines produced in these regions in the optical part of the spectrum and by means of photoionization models. We adapted the code HII-CHI-mistry to account for these abundances by implementing additional grids of models that assume a variable sulphur-to-oxygen abundance ratio, beyond the commonly assumed solar ratio. The addition of these models, and their use in a new iteration of the code allows us to use sulphur lines to precisely estimate the sulphur abundance, even in the absence of auroral lines. This approach aligns with the results from the direct method, and no additional assumptions about the ionization correction factor are needed, as the models directly predict the total sulphur abundance. We applied this new methodology to a large sample of star-forming regions from the MaNGA survey, and we explored the variation of the S/O ratio as a function of metallicity, making corrections for the significant contribution from diffuse ionized gas, which particularly affects the [S ii] emission. Our results indicate no significant deviations from the solar S/O value in the range 8.0 12+log(O/H) 8.7, where the bulk of the MaNGA sample stays, , but also with possible enhancements of sulphur production both at low and high metallicity regimes. The latter may be linked to the depletion of oxygen in the gas-phase due to its incorporation onto dust grains
@@ 1 Introduction
However, one of the main drawbacks of the optical regime to be used for the derivation of S abundance stays on the fact that a non-negligible fraction of the sulphur abundance can be found in higher stages of ionization, such as S+3, only measurable in the infra-red regime (e.g. at [S iv] 10.5 m) (Vermeij & van der Hulst, 2002). The acquisition of mid-IR data can thus be undertaken to have a more precise quantification of the total sulphur abundance, above all in nebula ionized by very hard incident radiation fields, such as Extreme Emission Line Galaxies (EELGs) or Active Galactic Nuclei (AGN). Therefore, if the observed spectrum is not complete, even if the direct method can be applied, additional assumptions can be made to calculate the total S abundance by means of an ionization correction factor (ICF) to account for the sulphur stages unseen in the optical or IR parts, contributing to enhance the uncertainty of the S abundance derivation (Stasińska, 1978; Pérez-Montero et al., 2006; Dors et al., 2016).
The emissivity of the auroral line from [S iii] at 6312 Å as an advantage in relation to the auroral lines emitted by other elements, such as [O iii[ 4363 Å, depends less on temperature, making it easier to detect at higher metallicity (e.g. Castellanos et al. 2002; Bresolin et al. 2005). Nevertheless, as the direct method for the determination of S abundances can hardly be applied in large samples composed of weak/distant objects, the estimation of S chemical abundance in the gas-phase must rely on the application of other techniques based on strong detected emission-lines. For instance, one of the most widely used parameters based uniquely on strong sulphur CELs is the parameter, defined by Vilchez & Esteban (1996) as:
@@
Another alternative to calculate abundances is the use of photoionization models, that when tailored to individual objects can account for the calculation of either ICF when auroral lines are detected, or the total S abundance from the unique use of strong lines, (e.g. Evans 1986; Garnett 1989; Pérez-Montero & Díaz 2007; Pérez-Montero et al. 2010), but only for a limited number of objects. On the other hand, the program HII-CHI-mistry (hereinafter HCm Pérez-Montero 2014)111All the different version of the program can be retrieved from the webpage http://home.iaa.es/~epm/HII-CHI-mistry.html uses another strategy to provide an estimate of the total elemental abundances based on a limited set of emission-lines with enough signal-to-noise ratio that can be applied to large number of objects.
The version of the code for the IR (HCm-IR, Fernández-Ontiveros et al. 2021; Pérez-Díaz et al. 2022; Pérez-Díaz et al. 2024a) is already prepared to calculate S abundances, taking advantage from the fact that S lines in the rest frame mid-IR have very low dependence on electron temperature, what enables the code to independently estimate O and S abundances. However, the situation in the optical regime is quite different, as S CELs are much more dependent on electron temperature. For this reason, in this paper we describe a new version of HCm for the optical rest frame spectral range to provide solutions for the S chemical abundance consistently with the direct method that can be applied even in absence of any auroral line, and we applied it to a large sample of star-forming regions in the MaNGA survey to shed some light on the behaviour of the S/O ratio with metallicity.
The paper is organized as follows: In Section 2, we describe the observational data compiled from several sources, including one sample with direct determinations of the sulphur abundance, used to evaluate the results of our code, and another larger sample to which the code was applied. In Section 3, we explain the method, detailing the grids of photoionization models and the HII-CHI-mistry code. In Section 4, we present and discuss our results. Finally, in Section 5, we summarize our findings and present our conclusions.
@@ 2 Data sample
For this work we compiled two main sources of observational information. The first sample has the aim of serving as a control sample for our methodology gathering optical spectral data, allowing a direct determination of their total sulphur abundance that can be used to evaluate the results of our code. The second one is a sample of pointings from the MaNGA survey selected as ionized mainly by star-formation. Although the quality of this data is good, the observations are not deep enough to derive total abundances based on the application of the direct method, relying on the detection of the weak auroral lines. Therefore, the application of other methods based on the use of strong emission lines is necessary.
@@ 2.1 Control sample of objects with direct sulphur abundance
We also took for our analysis the total S abundances calculated in this work, using ICFs based on the softness parameter or, in the case of observations only in the red part of the spectrum, based on Ar emission-lines. However, for O abundances, given that these are not provided in the same work for the whole sample, we recalculated them from the listed emission-lines, using pyneb (Luridiana et al., 2015), and adopting the empirical relation between oxygen and sulphur electron temperatures given by Hägele et al. (2006), following the procedure for the direct method described in Pérez-Montero (2017).
@@ 3 The models and the code
In this section we describe the computed photoionization model grids whose input conditions and predicted emission line fluxes are used to feed the code HCm to calculate chemical abundances and ionization parameter. . We also describe the new features of the code to derive S abundances, including the new considered observables and the calculation procedure.
@@ 3.1 Description of the photoionization models
Departing from the photoionization model grids already calculated for previous versions of the code HCm using Cloudy v. 17 (Ferland et al., 2017), we complemented these grids varying in each model different input values of the S/O ratio, including deviations from the solar ratio previously assumed.
The models calculated for star-forming regions used by the code cover a range for 12+log(O/H) from 6.9 to 9.1 in bins of 0.1 dex, for log(N/O) from -2.0 to 0.0 in bins of 0.125 dex, and for log 222From now on, we use the denomination to denote the ionization parameter used by HCm to underline the fact that the grid of models do not consider variations in the assumed incident SED or in the geometry, so the derived values depending on emission-line ratios only depend on according to the models. from -4.0 to -1.5 in bins of 0.25 dex. These models assume Popstar synthetic model atmospheres from Mollá et al. (2009) with an Initial Mass Function (IMF) from Chabrier (2003) at an age of 1 Myr. The gas distribution is assumed to have plane-parallel geometry, constant electron density of 100 cm-3, a filling factor of 0.1, and a standard galactic dust-to-gas mass ratio.
As a novelty, for each combination of O/H and ,, we calculated six additional models with different S/O ratio ranging from a deviation of -0.6 to 0.6 dex in relation to the solar value (log(S/O)⊙ = -1.57, Asplund et al. 2009) in bins of 0.2 dex. The rest of elements are still fixed to solar proportions according to the values given by Asplund et al. (2009), with the exception of N, which is fixed in these models with variable S/O according to the empirical relation found by Pérez-Montero (2014). Note that this is not incompatible with the variation of N/O explored in the models with solar S/O ratio, as the relative N abundance does not imply noticeable changes in the derivation of S abundance, once N/O has been calculated in previous steps, as described below, and the models with variable S/O ratio are not used to derive N abundances. These additional models with varied S/O were also incorporated to other libraries of model grids available for the code, including BPASS v.2.1 (Eldridge et al., 2017) models for its use for EELGs as described in Pérez-Montero et al. (2021). However, the inclusion of these new models and their use for the calculation of S/O has not been made for its use in the Narrow Line Regions (NLR) in AGN (Pérez-Montero et al., 2019a; Pérez-Díaz et al., 2024b) but it will be incorporated in the near future and discussed in more detail in forthcoming papers. Thus, for the above described types of ionizing source, mainly representing star-forming objects, the code incorporates two libraries listing the predicted emission line fluxes and the introduced initial conditions and abundances: one with solar S/O ratio for varying O/H, N/O, and ; and another one assuming an empirical relation between O/H and N/O with varying S/O, O/H, and .
@@ 3.2 Description of the code HII-CHI-mistry to derive sulphur abundances
The version 6.0 of HCm makes use of the model libraries described above, compiling the predicted emission-line fluxes and the model parameters required to calculate 12+log(O/H), log(N/O), log , and 12+log(S/H) with their corresponding uncertainties. The fact that our method finds these solutions from the total abundances taken as input in each model of the used grids is an advantage in relation to the direct method, as it implies that the calculation of ICFs is not necessary, step that it is often based on arbitrary assumptions made on biased samples or photoionization models that are not consistent with the rest of calculations. This advantage also extends to the fact that the code, contrary to the direct method, does not need to assume any arbitrary relation between the electron temperatures associated to the derivation of the corresponding chemical abundance of each ion, as the thermal structure is implicit in each model, and the ionic abundances are not necessary to calculate the elemental abundances in our method.
@@
The set of emission lines accepted by the code includes O ii] 3727 Å (sum of the doublet 3726+3729), [Ne iii] 3868 Å, [O iii] 4363 Å, [O iii] 4959,5007 Å (due to the theoretical ratio between these lines, the code admits only one of them), [N ii] 5755 Å, [S iii] 6312 Å, [N ii] 6584 Å, [S ii] 6716,6731 Å (provided separately or summed, but total flux must be specified), [O ii] 7325 Å (total flux of multiplet 7319-7330), and one or both of the doublet lines [S iii] 9069, 9532 Å. These fluxes must be reddening-corrected and given relative to H.
The code firstly performs an iteration through the entire grid of models to calculate N/O. This is derived as the weighted-mean of the N/O values of all models, using as weights of the value of the models. These are calculated as the quadratic difference between the observed and predicted values of certain observables. For this version, the parameter N2O2 (used as tracer of N/O by Pérez-Montero & Contini (2009) has priority, but N2S2 (Pérez-Montero & Contini, 2009) and O3N2 (Florido et al., 2022) can be also used. As in previous versions of the code, once N/O has been derived, its uncertainty is also estimated considering two independent sources of error: the standard deviation of the final -weighted N/O distribution, and the uncertainty associated to the given observational errors, calculated performing a Monte Carlo iteration perturbing the fiducial given input values with the uncertainties.
@@
Once N/O is fixed in the grid of models, the code performs a second iteration to provide an estimation for 12+log(O/H) and log , adopting the new denomination described in subsection 3.1. In this case, several observables can be used, depending on which of them are available as afuncion of the provided emission lines for each object. These are all described in Pérez-Montero (2014). Notice that, once N/O has been fixed in the previous step, observables based on [N ii] can be used without any previous arbitrary assumption on the O/H vs N/O relation.
@@
In this step, in addition to the observables based on the available optical CELs used by the code to build the resulting -weighted distributions to derive O/H and , we added those based on [S iii] emission lines that were not previously used in other versions of the code. Even though these lines present a certain dependence on the relative S abundance, some among them have a larger dependence on the total metallicity or on excitation.
@@
which has been mainly used as a proxy for the ionization parameter of the gas (Díaz, 1998). In right panel of Figure 1 we show the behaviour of the parameter as a function of log for different model sequences of the grid calculated for two values of 12+log(O/H) = 8.1 and 8.7, and showing also results for different S/O ratios around the solar value. As it can be seen, little variation is found as a function of the relative sulphur abundance and, consequently, this observable is only used in the second iteration of the code, which is aimed at the simultaneous search of both and O/H. Notice that this ratio has been also studied to derive the equivalent effective temperature of the incident field of radiation by Pérez-Montero et al. (2019b), as part of the input for the code HCm-Teff, so a more detailed determination of the ionization parameter should be made using it if a more precise grid attending to possible variations of the incident SED should be made. This partially explains the dispersion obtained for the studied objects in this plot, as the O2O3 ratio is also used for the derivation of . A more detailed discussion about the differences between and the ionization parameter will be made in subsection 4.2.
@@
The basis of this procedure is well justified by inspecting the behaviour of the model sequences shown in Figure 2, where is mainly driven by S/H, but with a large dependence on both O/H and . Therefore, an independent determination of these parameters prior to the derivation of S/H is necessary to supply a trustable and robust determination of the S/O ratio using the available S emission lines in the optical part of the spectrum.

### v2

@@ Abstract
The derivation of sulphur chemical abundances in the gas-phase of star-forming galaxies is explored in this work, using the emission lines produced in these regions in the optical part of the spectrum and by means of photoionization models. We adapted the code HII-CHI-mistry to account for these abundances by implementing additional grids of models that assume a variable sulphur-to-oxygen abundance ratio, beyond the commonly assumed solar value. The addition of these models, and their use in a new iteration of the code allows us to use sulphur lines to precisely estimate the sulphur abundance, even in the absence of auroral lines. This approach aligns with the results from the direct method, and no additional assumptions about the ionization correction factor are needed, as the models directly predict the total sulphur abundance. We applied this new methodology to a large sample of star-forming regions from the MaNGA survey, and we explored the variation of the S/O ratio as a function of metallicity, making corrections for the significant contribution from diffuse ionized gas, which particularly affects the [S ii] emission. Our results indicate no significant deviations from the solar S/O value in the range 8.0 12+log(O/H) 8.7, where the bulk of the MaNGA sample stays, but also with possible enhancements of sulphur at the high metallicity regime. This may be linked to the depletion of oxygen in the gas-phase due to its incorporation onto dust grains, as it remains when other metallicity indicators independent of this depletion, as S/H itself, are used instead.
@@ 1 Introduction
However, one of the main drawbacks of the optical regime to be used for the derivation of S abundance stays on the fact that a non-negligible fraction of the sulphur abundance can be found in higher stages of ionization, such as S+3, only measurable in the infra-red regime (e.g. at [S iv] 10.5 m) (Vermeij & van der Hulst, 2002). The acquisition of mid-IR data could thus be undertaken to have a more precise quantification of the total sulphur abundance, above all in nebula ionized by very hard incident radiation fields, such as Extreme Emission Line Galaxies (EELGs) or Active Galactic Nuclei (AGN). However, if the observed spectrum is not complete, even if the direct method can be applied, additional assumptions must be taken to calculate the total S abundance by means of an ionization correction factor (ICF) to account for the sulphur stages unseen in the optical or IR parts, contributing to enhance the uncertainty of the S abundance derivation (Stasińska, 1978; Pérez-Montero et al., 2006; Dors et al., 2016).
In addition, despite the fact that the emissivity of the auroral line from [S iii] at 6312 Å depends less on temperature than other optical auroral lines, such as [O iii] 4363 Å, , what makes easier to detect it at higher metallicity (e.g. Castellanos et al. 2002; Bresolin et al. 2005), the direct method for the determination of S abundances can hardly be applied in large samples composed of weak/distant objects. Therefore, the estimation of S chemical abundance in the gas-phase must rely on the application of other techniques based on strong detected emission-lines. For instance, one of the most widely used parameters based uniquely on strong sulphur CELs is the parameter, defined by Vilchez & Esteban (1996) as:
@@
Another alternative to calculate abundances is the use of photoionization models, that when tailored to individual objects can account for the calculation of either ICF when auroral lines are detected, or the total S abundance from the unique use of strong lines, (e.g. Evans 1986; Garnett 1989; Pérez-Montero & Díaz 2007; Pérez-Montero et al. 2010), but only for a limited number of objects. On the other hand, the program HII-CHI-mistry (hereinafter HCm Pérez-Montero 2014) uses another strategy to provide an estimate of the total elemental abundances based on a limited set of emission-lines with enough signal-to-noise ratio that can be applied to large number of objects.
There is a version of this code for the IR (HCm-IR, Fernández-Ontiveros et al. 2021; Pérez-Díaz et al. 2022; Pérez-Díaz et al. 2024a), that is already prepared to calculate S abundances, taking advantage from the fact that S lines in the rest frame mid-IR have very low dependence on electron temperature, what enables the code to independently estimate O and S abundances. However, the situation in the optical regime is quite different, as S CELs are much more dependent on electron temperature. For this reason, in this paper we describe a new version of HCm for the optical rest frame spectral range to provide solutions for the S chemical abundance consistently with the direct method that can be applied even in absence of any auroral line, and we applied it to a large sample of star-forming regions in the MaNGA survey to shed some light on the behaviour of the S/O ratio with metallicity.
The paper is organized as follows: In Section 2, we describe the observational data compiled from several sources, including one sample with direct determinations of the sulphur abundance, used to evaluate the results of our code, and another larger sample to which the code was applied. In Section 3, we explain the method, detailing the grids of photoionization models and the HCm code. In Section 4, we present and discuss our results. Finally, in Section 5, we summarize our findings and present our conclusions.
@@ 2 Data sample
For this work we compiled two main sources of observational information. The first sample has the aim of serving as a control sample for our methodology gathering optical spectral data, allowing a direct determination of their total sulphur abundance that can be used to evaluate the results of our code. The second one is a sample of pointings from the MaNGA survey selected as ionized mainly by star-formation. Although the quality of this MaNGA data is good, the observations are not deep enough to derive total abundances based on the application of the direct method, relying on the detection of the weak auroral lines. Therefore, the application of other methods exclusively based on the use of strong emission lines is necessary.
@@ 2.1 Control sample of objects with direct sulphur abundance
We also took for our analysis the total S abundances calculated in this work, using ICFs based on the softness parameter (Vílchez & Pagel, 1988) or, in the case of observations only in the red part of the spectrum, based on Ar emission-lines. However, for O abundances, given that these are not provided in the same work for the whole sample, we recalculated them from the listed emission-lines, using pyneb (Luridiana et al., 2015), and adopting the empirical relation between oxygen and sulphur electron temperatures given by Hägele et al. (2006), following the procedure for the direct method described in Pérez-Montero (2017).
@@ 3 The models and the code
In this section we describe the code HCm to calculate chemical abundances and the ionization parameter, , from a set of observed emission lines as compared to the predictions from a large grid of photoionization models. We focus on the new version of the code, version 6.0, capable of providing an estimate of the total sulphur abundance using optical [S ii] and [S iii] lines. We also describe the new considered observables based on these lines, the new added grid of models with a variable S/O ratio, and the calculation procedure.
@@ 3.1 Description of the photoionization models
In order to derive abundances, HCm operates using different grids of photoionization models defined by the user, but it incorporates by default libraries of emission line fluxes calculated with the code Cloudy v. 17 (Ferland et al., 2017). As in previous versions of the code (e.g. Pérez-Montero 2014; Pérez-Montero et al. 2021), for the derivation of O/H, N/O, and the models calculated for star-forming regions consider input conditions covering a range for 12+log(O/H) from 6.9 to 9.1 in bins of 0.1 dex, for log(N/O) from -2.0 to 0.0 in bins of 0.125 dex, and for log from -4.0 to -1.5 in bins of 0.25 dex. These models assume as incident radiation source Popstar synthetic model atmospheres from Mollá et al. (2009) with an Initial Mass Function (IMF) from Chabrier (2003) with an age of 1 Myr. The gas distribution is assumed to have plane-parallel geometry, constant electron density of 100 cm-3, a filling factor of 0.1, and a standard galactic dust-to-gas mass ratio.
In addition, as a novelty for this version and departing from the photoionization model grids already calculated for previous versions of the code, we also used Cloudy v. 17 to complement these grids varying in each model different input values of S/O, considering deviations from the solar ratio previously assumed in the original grid, but with the same values for the rest of input conditions. Then, for each combination of O/H and , we calculated six additional models with different S/O ratio ranging from a deviation of -0.6 to 0.6 dex in relation to the solar value (log(S/O)⊙ = -1.57, Asplund et al. 2009) in bins of 0.2 dex. The rest of elements were still fixed to solar proportions according to the values given by Asplund et al. (2009), with the exception of N, which was fixed in these additional models according to the empirical relation with O/H found by Pérez-Montero (2014). Note that this is not incompatible with the N/O variation considered in previous models, as this does not imply noticeable changes in the derivation of S/H. Overall, this additional grid with variable S/O implies a number of 1 518 models.
These additional models with varied S/O were also incorporated to other libraries of model grids available for the code, including BPASS v.2.1 (Eldridge et al., 2017) models for its use for EELGs as described in Pérez-Montero et al. (2021). However, the inclusion of these new models and their use for the calculation of S/O has not been made for its use in the Narrow Line Regions (NLR) in AGN (Pérez-Montero et al., 2019a; Pérez-Díaz et al., 2024b) but it will be incorporated in the near future and discussed in more detail in forthcoming papers. Thus, for the above described types of ionizing source, mainly representing star-forming objects, the code incorporates two libraries listing the predicted emission line fluxes and the introduced initial conditions and abundances: one with solar S/O ratio for varying O/H, N/O, and ; and another one assuming an empirical relation between O/H and N/O with varying S/O for each value of O/H, and .
@@ 3.2 Description of the code HII-CHI-mistry to derive sulphur abundances
HCm (Pérez-Montero, 2014) is a python-based script aimed at the derivation of certain chemical abundances and the ionization parameter in ionized gaseous nebulae, including star-forming regions, AGN, (Pérez-Montero et al., 2019a) or EELGs (Pérez-Montero et al., 2021). There are also specific versions able to deal with ultraviolet lines (Pérez-Montero & Amorín, 2017; Pérez-Montero et al., 2023) and the IR (Fernández-Ontiveros et al., 2021; Pérez-Díaz et al., 2022)111All the different version of the program can be retrieved from the webpage http://home.iaa.es/~epm/HII-CHI-mistry.html. This calculation is made from the comparison of the measured reddening-corrected fluxes of certain CELs with the prediction of large grids of photoionization models. The code basically estimates these quantities as the mean values of a -weighted distribution of all considered models, where the are the normalized quadratic difference between some observed and measured emission-line ratios sensitive to the searched quantities. The fact that our method finds these solutions as compared with the total abundances taken as input in each model of the used grids is an advantage in relation to the direct method, as it implies that the calculation of ICFs is not necessary, step that it is often based on arbitrary assumptions made on biased samples or photoionization models that are not necessaryly consistent with the rest of calculations. This advantage also extends to the fact that the code, contrary to the direct method, does not need to assume any arbitrary relation between the electron temperatures associated to the derivation of the corresponding chemical abundance of each ion, as the thermal structure is implicit in each model, and the ionic abundances are not necessary to calculate the elemental abundances in our method.
The version 6.0 of HCm for the optical makes use of the model libraries described above, compiling the predicted emission-line fluxes and the model parameters required to calculate 12+log(O/H), log(N/O), log , and 12+log(S/H) with their corresponding uncertainties, estimated considering two independent sources of error: the standard deviation of the final -weighted distribution, and the uncertainty associated to the given observational errors, calculated performing a Monte Carlo iteration perturbing the fiducial given input values with the uncertainties.
@@
The set of emission lines accepted by the code in this version includes O ii] 3727 Å (sum of the doublet 3726+3729), [Ne iii] 3868 Å, [O iii] 4363 Å, [O iii] 4959,5007 Å (due to the theoretical ratio between these lines, the code admits only one of them), [N ii] 5755 Å, [S iii] 6312 Å, [N ii] 6584 Å, [S ii] 6716,6731 Å (provided separately or summed, but total flux must be specified), [O ii] 7325 Å (total flux of multiplet 7319-7330), and one or both of the doublet lines [S iii] 9069, 9532 Å. These fluxes must be reddening-corrected and given relative to H.
The procedure and decission tree of the code has been designed as follows: The code firstly performs an iteration through the entire grid of models to calculate N/O, as the used observables to derive this ratio do not depend on . For this version, the parameter N2O2 (defined as log([N ii]/[O ii]), is used as tracer of N/O by Pérez-Montero & Contini (2009) with priority, but N2S2 (similar replacing [O ii] with [S ii])(Pérez-Montero & Contini, 2009) and O3N2 (Florido et al., 2022) can be also used.
@@
Once N/O is fixed in the grid of models, the code performs a second iteration to provide an estimation for 12+log(O/H) and log . In this case, several observables can be used, depending on which of them are available as a function of the provided emission lines for each object. These are all described in Pérez-Montero (2014). Notice that, once N/O has been fixed in the previous step, the code constrains the grid to the found N/O values, and the observables based on [N ii] can then be used without any previous arbitrary assumption on the O/H vs N/O relation. Nevertheless, in case that a previous N/O estimation has not been possible, an empirical relation between O/H and N/O is adopted instead, although this may imply additional sources of uncertainty if the studied object does not follow it. In addition, the code also constrains the grid in case that any auroral to nebular emission line ratio (e.g. [O iii] 4363/5007 Å) is not given, following an empirical relation between O/H and , that will be discussed in next section.
@@
For this version of the code dealing with S emission lines, in this step, in addition to the observables based on the available optical CELs to build the resulting -weighted distributions to derive O/H and , we added those based on [S iii] emission lines that were not previously used in other versions of the code. Even though these lines present a certain dependence on the relative S abundance, some among them have a larger dependence on the total metallicity or on excitation.
@@
which has been mainly used as a proxy for the ionization parameter of the gas (Díaz, 1998). In right panel of Figure 1 we show the behaviour of the parameter as a function of log for different model sequences of the grid calculated for two values of 12+log(O/H) = 8.1 and 8.7, and showing also results for different S/O ratios around the solar value. As it can be seen, little variation is found as a function of the relative sulphur abundance and, consequently, this observable is only used in the second iteration of the code, which is aimed at the simultaneous search of both and O/H. Notice that this ratio has been also studied to derive the equivalent effective temperature of the incident field of radiation by Pérez-Montero et al. (2019b), as part of the input for the code HCm-Teff, so a more detailed determination of the ionization parameter should be made using it if a more precise grid attending to possible variations of the incident SED should be made. This partially explains the dispersion obtained for the studied objects in this plot, as the O2O3 ratio is also used for the derivation of .
@@
The basis of this procedure is well justified by inspecting the behaviour of the model sequences shown in Figure 2, where is mainly driven by S/H, but with a large dependence on both O/H and . Therefore, an independent determination of these parameters prior to the derivation of S/H is necessary to supply a trustable and robust determination of the S/H ratio using the available S emission lines in the optical part of the spectrum.
As a first sanity check of the results obtained by the code, we performed a test consisting of using as input the emission-line fluxes predicted by the models to find out if the code obtains the same chemical abundances used by the models. The results of this operation are shown in Table LABEL:test, for which we give the mean offsets and the standard deviations between the calculated O/H and S/H and the corresponding input values using the popstar models. We explored different scenarios including calculations made using: all possible lines, both nebular and auroral; using only strong nebular lines; both auroral and nebular but only in the red part of the spectrum (i.e. from [S iii] 6312Å); and, finally, only strong lines in this same regime. We provide results when no error is assigned to the input fluxes and, between parenthesis, the result when a 10% error is considered.
| Case | 12+log(O/H) | 12+log(S/H) | ||||
| Mean error | Mean | St.dev. | Mean error | Mean | St.dev. | |
| All222Only models which predict a flux for the auroral lines | 0.000 (0.093) | +0.000 (+0.018) | 0.003 (0.058) | 0.027 (0.040) | -0.011 (+0.007) | 0.095 (0.113) |
| Strong lines333Only models in the adopted O/H- relation | 0.000 (0.059) | +0.001 (+0.012) | 0.031 (0.069) | 0.016 (0.021) | -0.003 (+0.018) | 0.138 (0.168) |
| Red rangea | 0.002 (0.192) | +0.005 (-0.017) | 0.064 (0.258) | 0.025 (0.039) | -0.003 (+0.005) | 0.099 (0.145) |
| Red strong linesb | 0.000 (0.128) | -0.000 (+0.014) | 0.007 (0.145) | 0.015 (0.024) | -0.013 (+0.020) | 0.132 (0.210) |
The resulting values obtained from the comparison between the results from HCm with the values in the models shows that we can recover the chemical abundances using only emission-lines with a high precision, although this can be altered when the uncertainty in the measurement of the corresponding lines are also considered. This uncertainty can be due to very different causes, including extinction or aperture factors, but their impact on the results of our method should be considered by the user when include certain emission lines.